OpenStudy (anonymous):

x^2+y^2=9 16x^2-4y^2=64

OpenStudy (anonymous):

If you have an equation that can be written as x^2/a^2 + y^2/b^2 = 1 (a>=b>0) it is an ellipse (circle is a special case of an ellipse) centred on the origin. If you have an equation that can be written as x^2/a^2 - y^2/b^2 = 1 (a,b>0) it is a hyperbola centred on the origin.

OpenStudy (anonymous):

solve for real solutions

OpenStudy (anonymous):

Ah, you didn't say that:-) If there are real solutions there must be 4 of them (where the arms of the hyperbola cross the ellipse.

OpenStudy (anonymous):

what are they though

OpenStudy (anonymous):

Let's see, if I work out 1 of them that will give me the other three by symmetry...

OpenStudy (anonymous):

ok so whata are all 4 ?

OpenStudy (anonymous):

Hmm...multiply that first one by 16 and subtract the second one from it gives 20y^2 = 80 or y= 2.....

OpenStudy (anonymous):

Well, plus or minus 2....

OpenStudy (anonymous):

so what are all 4

OpenStudy (anonymous):

can you type them all out for me

OpenStudy (anonymous):

I will, just on the phone...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Huh, just noticed I can divide that second one by 4...duh.

OpenStudy (anonymous):

How about (root5, plus or minus 2) and (-root5, plus or minus 2)?

OpenStudy (anonymous):

yeah thats right thank you

OpenStudy (anonymous):

u r welcome:-)

jhonyy9 (jhonyy9):

x*2 +y*2 =9 16x*2 -4y*2 =64 /4 --- 4x*2 -y*2 =16 --- y*2 = 4x*2 -16 x*2 +4x*2 -16 =9 5x*2 =25 x*2 =5 x_1,_2 = +/- radical 5